OpenStudy (anonymous):

the speed of a passenger train is 12mph faster than of a frieght train.the passenger train travels 330 miles in the same time it takes the frieght train to travel 270 miles find the speed of each train

OpenStudy (anonymous):

let x be the speed of the slower train 11/9*x = x + 12 9x + 108 = 11 x 2x = 108 x = 54 x+12 = 66